OpenStudy (anonymous):

there are 75 runners to start a race. in the first half of the race2/5 of them dropped out. in the second half of the race, 3/5 of the remaining runners dropped out. How many runners finished the rce?

OpenStudy (anonymous):

its 27

OpenStudy (epoweritheta):

after first half # runers = 3/5*75=45 , after 2nd half # runers=45*2/5=18

OpenStudy (anonymous):

18 is the correct answer i forgot some math

OpenStudy (epoweritheta):

CelesteSelene13 is wrong

OpenStudy (anonymous):

75*2/5=30 75-30=45 because 30 people dropped out 45*3/5=27 45-27=18 27 people dropped out in the second part so 18 people finished the race.